OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) =4/x. and g(x) =4 /x

OpenStudy (anonymous):

okay let's put y=g(x) then f(g(x))=f(y)=4/y now you replace y by g(x) f(g(x))=4/g(x) now you replace g(x) by its expression f(g(x))=4/(4/x)=4*(x/4)=x

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

the same thing for g(f(x)) you're welcome :)